OpenStudy (anonymous):

In a geometric sequence where r > 1, the terms always increase. True or False??

ganeshie8 (ganeshie8):

say r = 1.1 a = 1 then the terms will be \(1, 1\times 1.1, 1\times 1.1^2, ..... \)

ganeshie8 (ganeshie8):

is \(1\times 1.1 > 1\) ?

OpenStudy (anonymous):

yes because 1.1 > 1, so then that would be true?

ganeshie8 (ganeshie8):

Yup !

OpenStudy (anonymous):

thank you so much!

ganeshie8 (ganeshie8):

wait a second, if a < 1, then the terms may decrease as well

OpenStudy (anonymous):

it was true :)

ganeshie8 (ganeshie8):

but i feel the quesiton is asking u the case when a is positive only. so the answer should be true only

ganeshie8 (ganeshie8):

cool :)

OpenStudy (anonymous):

thank you so much!